Superior Micro Products uses the weighted-average method in its process costing system. During January, the Delta Assembly Department completed its processing of 25,500 units and transferred them to the next department. The cost of beginning work in process inventory and the costs added during January amounted to $705,740 in total. The ending work in process inventory in January consisted of 3,100 units, which were 80% complete with respect to materials and 60% complete with respect to labor and overhead. The costs per equivalent unit for the month were as follows: Materials Labor Overhead Cost per equivalent unit $ 13.00 $ 4.90 $ 7.60 Required: 1. Compute the equivalent units of materials, labor, and overhead in the ending work in process inventory for the month. 2. Compute the cost of ending work in process inventory for materials, labor, overhead, and in total for January. 3. Compute the cost of the units transferred to the next department for materials, labor, overhead, and in total for January. 4. Prepare a cost reconciliation for January. (Note: You will not be able to break the cost to be accounted for into the cost of beginning work in process inventory and costs added during the month.)
Answer #1

1) Equivalent unit

Material Labor Overhead
equivalent unit for ending WIP 3100*80% = 2480 3100*60% = 1860 1860

2) Total cost of ending WIP

Material Labor Overhead Total
Cost of ending WIP 2480*13 = 32240 1860*4.90 = 9114 1860*7.60 = 14136 55490

3) Total cost of Unit transferred out

Material Labor Overhead Total
Cost of Unit transferred out 331500 124950 193800 650250

4) Cost reconciliation

Cost to be accounted for 705740
Cost accounted as
Cost of unit transferred out 650250
Cost of ending WIP 55490
Total cost accounted as 705740